# STAPO: Stabilizing Reinforcement Learning for LLMs by Silencing Rare Spurious Tokens

### v1

@@ Abstract
Reinforcement Learning (RL) has significantly improved large language model reasoning, but existing RL fine-tuning methods rely heavily on heuristic techniques such as entropy regularization and reweighting to maintain stability. In practice, they often experience late-stage performance collapse, leading to degraded reasoning quality and unstable training. We derive that the magnitude of token-wise policy gradients in RL is negatively correlated with token probability and local policy entropy. Building on this result, we prove that training instability is driven by a tiny fraction of tokens, approximately 0.01%, which we term spurious tokens. When such tokens appear in correct responses, they contribute little to the reasoning outcome but inherit the full sequence-level reward, leading to abnormally amplified gradient updates. Motivated by this observation, we propose Spurious-Token-Aware Policy Optimization (STAPO) for large-scale model refining, which selectively masks such updates and renormalizes the loss over valid tokens. Across six mathematical reasoning benchmarks using Qwen 1.7B, 8B, and 14B base models, STAPO consistently demonstrates superior entropy stability and achieves an average performance improvement of 7.13% over GRPO, 20-Entropy and JustRL.
@@
Figure 1: (a) Conceptual analogy: We argue that spurious tokens, which are rare and uninformative tokens within otherwise correct responses that receive disproportionately large gradient updates, can harm training stability, analogous to a dissonant vocalist disrupting the harmony of a performance. (b) By masking this negligible fraction (near $0.01\%$) of spurious tokens during the RL process of Qwen3-8B-Base, STAPO approaches the Pareto frontier of performance (AIME24 Acc) and entropy stability, compared to GRPO [1], 20-Entropy [2], and JustRL [3].
@@ 1 Introduction
Recent large language models (LLMs), including OpenAI-o1 [4], DeepSeek-R1 [5], and Qwen3 [6], have demonstrated remarkable capabilities in complex reasoning domains such as mathematics and programming. Central to this success is Reinforcement Learning (RL), which optimizes directly for outcome-level correctness and has been empirically linked to the emergence of advanced, long-horizon reasoning behaviors [7]. However, training stability remains a critical bottleneck in RL. Models often suffer from catastrophic degradation, abruptly collapsing from coherent reasoning chains into shallow, repetitive, or nonsensical patterns [8].
Prior work has approached this instability from two perspectives: entropy regulation and gradient modulation. On the entropy front, methods such as selective regularization [2], sample augmentation [9], and clipping modifications [10, 11, 12] attempt to arrest entropy collapse. However, these interventions often precipitate the opposite mode, inducing oscillatory or excessive entropy that degrades reasoning coherence [13]. Turning to the gradient modulation perspective, the disproportionate impact of rare tokens on parameter updates has motivated efforts to suppress unstable gradients via advantage reweighting [14], probability shaping [15], or spark-preserving regularization [16]. Nevertheless, existing approaches lack a fine-grained diagnosis of token instability. By overlooking the interplay between token probability and policy entropy, they fail to distinguish varying degrees of model confidence within low-probability regions, conflating valuable exploration with detrimental noise and leaving the structural sources of instability.
@@
  • We derive that the magnitude of token-wise policy gradients in RL is negatively correlated with token probability and local policy entropy. We further establish a unified framework for analyzing training stability in the joint space of policy entropy, token probability, and advantage sign.
  • We identify a pathological class of low-probability, low-entropy tokens with positive advantage, termed spurious tokens, whose amplified gradients destabilize optimization. To address this, we propose Spurious-Token-Aware Policy Optimization (STAPO), which masks these tokens (approximately $0.01\%$) and renormalizes the loss over valid tokens.
  • Across six mathematical reasoning benchmarks and three model scales, STAPO substantially stabilizes policy entropy and consistently improves reasoning performance under diverse evaluation settings, demonstrating its robustness and effectiveness in stabilizing RL training.
@@ 2.2 Group Relative Policy Optimization (GRPO)
We briefly review Group Relative Policy Optimization (GRPO) [17], which estimates advantages without an explicit value function. For each prompt $\bm{x}$, GRPO samples a group of $G$ output sequences $\{\bm{y}_{1},\dots,\bm{y}_{G}\}$ from a reference policy $\pi_{\theta_{\mathrm{old}}}$. The optimization objective is defined as the average clipped surrogate loss over the group:
@@ 2.3 Clip-Higher and Token Normalization
Building upon the GRPO objective, DAPO [10] proposes four additional training tricks, among which token-level normalization and an asymmetric clip-higher mechanism are shown to be particularly effective for stabilizing RL training [18, 11, 12]. JustRL [3] adopts these two components and achieves strong empirical performance. The corresponding objective is written as
@@ 3.1 Token-Level Optimization Mechanisms in RL Training
Understanding how RL-style optimization updates reasoning models is central to analyzing training stability. Prior work has examined entropy dynamics [19, 12] or gradient behavior separately [14, 16]. In contrast, we provide a unified analysis that jointly characterizes gradient scaling and entropy evolution under group-style updates.
@@
3.2 Spurious Tokens as a Destructive Update Mode
@@ 3.3 Spurious-Token-Aware Policy Optimization (STAPO)
$\mathcal{J}_{\text{STAPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}\bm{x}\sim\mathcal{D}\\ \{\bm{y}_{i}\}_{i=1}^{G}\sim\pi_{\mathrm{old}}\end{subarray}}\left[\frac{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sum_{i=1}^{G}\sum_{t=1}^{|\bm{y}_{i}|}\mathbb{I}^{\text{S2T}}_{i,t}}}\sum_{i=1}^{G}\sum_{t=1}^{|\bm{y}_{i}|}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{I}^{\text{S2T}}_{i,t}}\cdot\min\left(\rho_{i,t}\hat{A}_{i},\text{clip}(\rho_{i,t},1-\epsilon,1+\epsilon)\hat{A}_{i}\right)\right].$ (9)
@@
Comparing Eq. (9) with the standard DAPO objective, there are two main differences highlighted in red. Firstly, STAPO selectively masks the loss calculation for spurious tokens. Furthermore, the normalization term in Eq. (9) is adjusted to account only for the remaining valid tokens. This design effectively prevents the misleading update due to the stochastic sampling process, thereby stabilizing the RL alignment process. The overall algorithm is summarized in Algorithm 1.
@@ 4 Related Work
Reinforcement Learning for LLMs. In contrast to traditional reinforcement learning paradigms applied to control and decision-making problems [21, 22, 23], reinforcement learning from human feedback (RLHF) has emerged as a predominant approach for aligning LLMs with human preferences and diverse downstream objectives [4, 5]. While early approaches relied on on-policy optimization such as PPO [24], recent work has shifted toward more efficient preference-based DPO [25], which avoids explicit reward modeling and online rollouts. The focus has further expanded from general alignment to improving reasoning capabilities, motivating RL and training schemes such as GRPO [17]. In this context, a range of policy optimization variants, including DAPO [10], GSPO [18], SAPO [11], and other related methods [26, 27, 9], have been developed to enhance optimization stability, sample efficiency, and scalability for reasoning-oriented language models.
@@
Entropy Instability in RL. A persistent challenge in training reasoning-oriented language models with RL is the rapid collapse of policy entropy during early optimization. Prior work mitigates this issue through interventions such as selectively regularizing high-entropy tokens [2], increasing the proportion of entropy-enhancing samples [12], and modifying clipping strategies in policy optimization [10, 28, 11]. However, these methods often introduce the opposite failure mode, where entropy grows excessively or becomes unstable, degrading reasoning coherence and leading to repetitive or unstructured outputs. Although some studies analyze entropy dynamics during RL training [19, 20] and others explicitly enforce entropy stability as an optimization objective [13], existing approaches largely treat entropy as a surface-level training signal rather than addressing the underlying sources of instability.
@@
Gradient Domination by Low-Probability Tokens. A key microscopic source of instability stems from the disproportionate gradient influence of low-probability tokens. As proved by [14], rare tokens can generate excessively large updates, allowing a small subset of unstable predictions to dominate optimization, which harms fine-tuning stability. Recent work addresses this through probability-aware modulation. Yang et al. [14] propose Advantage Reweighting and Low-Probability Token Isolation to attenuate or asynchronously update rare tokens, while Token-Regulated GRPO (TR-GRPO) [15], using probability shaping to suppress sharp directions induced by low-confidence predictions. To avoid suppressing informative exploratory signals, Low-Probability Regularization (Lp-Reg) [16], which filters noise while preserving meaningful rare tokens, and GTPO [29] mitigates conflicting gradients arising from such tokens across reward trajectories. However, these approaches largely rely on scalar probability thresholds, lacking a joint, fine-grained treatment of token-level confidence and probability, and thus failing to distinguish useful exploration from aleatoric noise under local model calibration.
@@ 5.1 Settings
Benchmarks. We evaluate our models on six widely adopted and challenging mathematical reasoning benchmarks: AIME24 [32], AIME25 [33], AMC23 [32] , MATH500 [34], Minerva [35], and OlympiadBench [36]. We generate $N$ independent responses per problem ($N=4$ for MATH500, Minerva, and OlympiadBench; $N=32$ for others) across two decoding configurations: temperature=1.0, top-p=1.0 and temperature=0.7, top-p=0.9, with a maximum length of 20,480 tokens. Results are reported as the average accuracy. To ensure evaluation rigour, we employ CompassVerifier-3B [37], a lightweight LLM verifier, to rectify misclassifications from the rule-based verification.
@@
Figure 3: Training Results across Different Model Scales. Each row presents the training dynamics for a specific model size (1.7B, 8B, and 14B). Notably, our proposed STAPO method achieves superior performance while maintaining stable policy entropy across all model sizes.
@@ 5.2.2 Performance under Training-Aligned Settings
Table 2 summarizes the main results. We first focus on the Training-Aligned evaluation setting (temperature=1.0, top-p=1.0), which reflects the model’s intrinsic capability without the aid of greedy decoding strategies. Under this setting, STAPO demonstrates consistent scalability and superior performance across all model sizes. For the 1.7B model, STAPO achieves a substantial lead, surpassing the strongest baseline (20-Entropy) by a relative margin of 13.50% in average accuracy. At the 8B scale, STAPO continues to outperform all competing methods, maintaining its position as the state-of-the-art. The efficacy of our method is further validated at the 14B scale, where STAPO extends its lead, achieving a relative improvement of approximately 5.94% over the best-performing baseline. These results collectively demonstrate that STAPO effectively unlocks the reasoning potential of base models across varying scales.
5.2.3 Performance Across Evaluation Settings
Recent works adopt different evaluation configurations [10, 2, 3, 12]. To ensure a fair comparison, we additionally evaluate STAPO under the JustRL setting (shown in parentheses/gray in Table 2). The results indicate that STAPO remains the top-performing method across different model scales under this setting. However, the performance gap between STAPO and the baselines becomes slightly narrower compared to the training-aligned configuration. We attribute this effect to the evaluation protocol: baselines with unstable, high-entropy distributions benefit more from constrained decoding, which suppresses low-probability “tail” generations. In contrast, STAPO learns a more intrinsically stable distribution and therefore relies less on decoding heuristics. Overall, STAPO achieves state-of-the-art performance under both configurations, demonstrating its robustness and stability.
Table 2: Main Results on Six Mathematical Reasoning Benchmarks across Three Model Scales. Each cell reports performance under the training-aligned configuration (temperature = 1.0, top-p = 1.0) and the JustRL [3] evaluation setting (temperature = 0.7, top-p = 0.9, shown in gray). For each configuration, the best result is highlighted in bold (training-aligned setting) and in dark gray (JustRL setting), respectively.
@@ 5.2.3 Performance Across Evaluation Settings
(a) Sensitivity to Probability Threshold ($\tau_{p}$)
(b) Sensitivity to Entropy Threshold ($\tau_{h}$)
Figure 4: Sensitivity Analysis of Probability Threshold $\tau_{p}$ and Entropy Threshold $\tau_{h}$ for Qwen3-1.7B Base. Performance is reported as Acc (avg@32) on AIME24 and AIME25 benchmarks.
@@ 5.3 Hyperparameter Sensitivity
(a) Average Acc (avg@32) across six benchmarks with temperature=1.0, top-p=1.0.
(b) Average Acc (avg@32) across six benchmarks with temperature=0.7, top-p=0.9.
Figure 5: Ablation Results of Different Masking Strategies across Three Model Scales.

### v2

@@ Abstract
Reinforcement Learning (RL) has significantly improved large language model reasoning, but existing RL fine-tuning methods rely heavily on heuristic techniques such as entropy regularization and reweighting to maintain stability. In practice, they often suffer from late-stage performance collapse, leading to degraded reasoning quality and unstable training. Our analysis shows that the magnitude of token-wise policy gradients in RL is negatively correlated with token probability and local policy entropy. We find that training instability can be caused by a tiny fraction of tokens, approximately 0.01%, which we term spurious tokens. When such tokens appear in correct responses, they contribute little to the reasoning outcome but inherit the full sequence-level reward, leading to abnormally amplified gradient updates. To mitigate this instability, we design S2T (silencing spurious tokens) mechanism to efficiently identify spurious tokens through characteristic signals with low probability, low entropy, and positive advantage, and then to suppress their gradient perturbations during optimization. Incorporating this mechanism into a group-based objective, we propose Spurious-Token-Aware Policy Optimization (STAPO), which promotes stable and effective large-scale model refinement. Across six mathematical reasoning benchmarks using Qwen 1.7B, 8B, and 14B base models, STAPO consistently demonstrates superior entropy stability and achieves an average performance improvement of 7.13% ($\rho_{\mathrm{T}}$=1.0, top-p=1.0) and 3.69% ($\rho_{\mathrm{T}}$=0.7, top-p=0.9) over GRPO, 20-Entropy and JustRL.
@@
S. E. Li and J. Duan with email .
Figure 1: Core Idea. (a) Conceptual analogy: We argue that spurious tokens, which are rare and uninformative tokens within otherwise correct responses that receive disproportionately large gradient updates, can harm training stability, analogous to a dissonant vocalist disrupting the harmony of a performance. (b) By masking this negligible fraction (near $0.01\%$) of spurious tokens during the RL process of Qwen3-8B-Base, STAPO approaches the Pareto frontier of performance (AIME24 Acc) and entropy stability, compared to GRPO, 20-Entropy, and JustRL.
@@ 1 Introduction
Recent large language models (LLMs), including OpenAI-o1, DeepSeek-R1, and Qwen3, have demonstrated remarkable capabilities in complex reasoning domains such as mathematics and programming. Central to this success is Reinforcement Learning (RL) [1], which optimizes directly for outcome-level correctness and has been empirically linked to the emergence of advanced, long-horizon reasoning behaviors [2]. However, training stability remains a critical bottleneck in RL. Models often suffer from catastrophic degradation, abruptly collapsing from coherent reasoning chains into shallow, repetitive, or nonsensical patterns [3].
Prior work has approached this instability from two perspectives: entropy regulation and gradient modulation. On the entropy front, methods such as selective regularization [4], sample augmentation [5], and clipping modifications [6, 7] attempt to arrest entropy collapse. However, these interventions often precipitate the opposite mode, inducing oscillatory or excessive entropy that degrades reasoning coherence [8]. Turning to the gradient modulation perspective, the disproportionate impact of rare tokens on parameter updates has motivated efforts to suppress unstable gradients via advantage reweighting [9], probability shaping [10], or spark-preserving regularization [11]. Nevertheless, existing approaches lack a fine-grained diagnosis of token instability. By overlooking the interplay between token probability and policy entropy, they fail to distinguish varying degrees of model confidence within low-probability regions, conflating valuable exploration with detrimental noise and leaving the structural sources of instability.
@@
  • Our analysis shows that the magnitude of token-wise policy gradients in RL is negatively correlated with token probability and local policy entropy. We further establish a unified framework for analyzing training stability in the joint space of policy entropy, token probability, and advantage sign.
  • We identify spurious tokens, a pathological class of tokens that appear in correct responses but contribute little to the reasoning process, whose disproportionately amplified gradients distort optimization and destabilize training. To address this issue, we introduce the S2T (silencing spurious tokens) mechanism, which efficiently detects spurious tokens characterized by low probability, low entropy, and positive advantage, and then suppresses their gradient perturbations during optimization. By incorporating S2T into a group-based objective, we propose STAPO, which effectively stabilizes large-scale model refinement.
  • Across six mathematical reasoning benchmarks (AIME24, AIME25, AMC23, MATH500, Minerva, and OlympiadBench) and three model scales (Qwen 1.7B, 8B, and 14B), STAPO substantially stabilizes policy entropy and consistently improves reasoning accuracy under diverse evaluation settings, demonstrating its SOTA performance in large-scale model refining tasks.
@@ 2.2 Group Relative Policy Optimization (GRPO)
We briefly review Group Relative Policy Optimization (GRPO) [12], which estimates advantages without an explicit value function. For each prompt $\bm{x}$, GRPO samples a group of $G$ output sequences $\{\bm{y}_{1},\dots,\bm{y}_{G}\}$ from a behaviour policy $\pi_{\theta_{\mathrm{old}}}$. The optimization objective, which eliminates the directly imposed KL penalty term, is defined as the average clipped surrogate loss over the group
@@ 2.3 Clip-Higher and Token Normalization
Starting from GRPO design, a few large-scale RL algorithms were proposed last year. DAPO [6] proposes four additional training tricks, among which token-level normalization and an asymmetric clip-higher mechanism are shown to be particularly effective for stabilizing RL training. JustRL [13] adopts these two components and achieves strong empirical performance. The corresponding objective is written as
@@ 3.1 Token-Level Optimization Mechanisms in RL Training
Understanding how RL-style optimization updates reasoning models is central to analyzing training stability. Prior work has examined entropy dynamics [14] or gradient behavior separately [9]. In contrast, we provide a unified analysis that jointly characterizes gradient scaling and entropy evolution under group-style updates.
@@
3.2 Spurious Token as Destructive Update Mode
@@ 3.3 Spurious-Token-Aware Policy Optimization (STAPO)
$\mathcal{J}_{\text{STAPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}\bm{x}\sim\mathcal{D}\\ \{\bm{y}_{i}\}_{i=1}^{G}\sim\pi_{\mathrm{old}}\end{subarray}}\left[\frac{1}{{\sum_{i=1}^{G}\sum_{t=1}^{|\bm{y}_{i}|}\mathbb{I}^{\text{S2T}}_{i,t}}}\sum_{i=1}^{G}\sum_{t=1}^{|\bm{y}_{i}|}{\mathbb{I}^{\text{S2T}}_{i,t}}\cdot\min\left(\rho_{i,t}(\theta)\hat{A}_{i},\text{clip}(\rho_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i}\right)\right].$ (9)
@@
Comparing Eq. (4) with the standard DAPO objective, there are two main differences. Firstly, STAPO selectively masks the loss calculation for spurious tokens. Furthermore, the normalization term in Eq. (9) is adjusted to account only for the remaining valid tokens. This design effectively prevents the misleading update due to the stochastic sampling process, thereby stabilizing the RL alignment process. The overall algorithm is summarized in Algorithm 1.
@@ 4 Related Work
Reinforcement Learning for LLMs. In the era of small model training, various effective RL algorithms have been proposed, for example, TD3 [18], TRPO [19], DSAC [20], and BOOM [21]. Recently, RL from human feedback (RLHF) has emerged as a predominant approach for aligning LLMs with human preferences and diverse downstream objectives [22, 23]. While early approaches relied on on-policy optimization such as PPO [24], recent work has shifted toward more efficient preference-based DPO [25], which avoids explicit reward modeling and online rollouts. The focus has further expanded from general alignment to improving reasoning capabilities, motivating RL and training schemes such as GRPO [12]. In this context, a range of policy optimization variants, including DAPO [6], GSPO [26], SAPO [7], and other related methods [27, 28, 5], have been developed to enhance optimization stability, sample efficiency, and scalability for reasoning-oriented language models.
@@
Entropy Instability in RL. Entropy is a central issue in RL, as it directly affects exploration and training stability [29, 30]. In reasoning-oriented language models trained with RL, a persistent challenge is the rapid collapse of policy entropy during the early stages of optimization, which often leads to premature convergence and degraded performance. Prior work mitigates this issue through interventions such as selectively regularizing high-entropy tokens [4], increasing the proportion of entropy-enhancing samples [15], and modifying clipping strategies in policy optimization [6, 31, 7]. However, these methods often introduce the opposite failure mode, where entropy grows excessively or becomes unstable, degrading reasoning coherence and leading to repetitive or unstructured outputs. Although some studies analyze entropy dynamics during RL training [14, 16] and others explicitly enforce entropy stability as an optimization objective [8], existing approaches largely treat entropy as a surface-level training signal rather than addressing the underlying sources of instability.
@@
Gradient Domination by Low-Probability Tokens. Gradients are a key factor in the stability of RL, and various forms of policy gradients have been proposed, for example, bicriteria policy gradient [32] and continuous-time policy gradient [33]. A key microscopic source of instability stems from the disproportionate gradient influence of low-probability tokens. As proved by [9], rare tokens can generate excessively large updates, allowing a small subset of unstable predictions to dominate optimization, which harms fine-tuning stability. Recent work addresses this through probability-aware modulation. To avoid suppressing informative exploratory signals, Low-Probability Regularization (Lp-Reg) [11] filters noise while preserving meaningful rare tokens. However, these approaches largely rely on scalar probability thresholds, lacking a joint, fine-grained treatment of token-level confidence and probability, and thus failing to distinguish useful exploration from aleatoric noise under local model calibration.
@@ 5.1 Settings
Benchmarks. We evaluate our models on six widely adopted and challenging mathematical reasoning benchmarks: AIME24 [35], AIME25 [36], AMC23 [35] , MATH500 [37], Minerva [38], and OlympiadBench [39]. We generate $N$ independent responses per problem ($N=4$ for MATH500, Minerva, and OlympiadBench; $N=32$ for others) across two decoding configurations: temperature $\rho_{\mathrm{T}}$=1.0, top-p=1.0 and temperature $\rho_{\mathrm{T}}$=0.7, top-p=0.9, with a maximum length of 20,480 tokens. Results are reported as the average accuracy. To ensure evaluation rigour, we employ CompassVerifier-3B [40], a lightweight LLM verifier, to rectify misclassifications from the rule-based verification.
@@
Figure 3: Training Results across Different Models. Each row presents the training dynamics for a specific model size (1.7B, 8B, and 14B). Notably, STAPO achieves superior performance while maintaining stable policy entropy across all model sizes.
@@ 5.2.2 Performance under Training-Aligned Settings
Table 2 summarizes the main results. We first focus on the Training-Aligned evaluation setting (temperature $\rho_{\mathrm{T}}$=1.0, top-p=1.0), which reflects the model’s intrinsic capability without the aid of greedy decoding strategies. Under this setting, STAPO demonstrates consistent scalability and superior performance across all model sizes. For the 1.7B model, STAPO achieves a substantial lead, surpassing the strongest baseline (20-Entropy) by a relative margin of 13.50% in average accuracy. At the 8B scale, STAPO continues to outperform all competing methods, maintaining its position as the state-of-the-art. The efficacy of our method is further validated at the 14B scale, where STAPO extends its lead, achieving a relative improvement of approximately 5.94% over the best-performing baseline. These results collectively demonstrate that STAPO effectively unlocks the reasoning potential of base models across varying scales.
Table 2: Main Results on Six Benchmarks across Three Models. Each cell reports performance under the training-aligned configuration (temperature $\rho_{\mathrm{T}}$= 1.0, top-p = 1.0) and the JustRL [13] evaluation setting (temperature $\rho_{\mathrm{T}}$= 0.7, top-p = 0.9, shown in gray). For each configuration, the best result is highlighted in bold (training-aligned setting) and in dark gray (JustRL setting), respectively.
@@
5.2.3 Performance Across Evaluation Settings
Recent works adopt different evaluation configurations [6, 4, 13, 15]. To ensure a fair comparison, we additionally evaluate STAPO under the JustRL setting (shown in parentheses/gray in Table 2). The results indicate that STAPO remains the top-performing method across different model scales under this setting. However, the performance gap between STAPO and the baselines becomes slightly narrower compared to the training-aligned configuration. We attribute this effect to the evaluation protocol: baselines with unstable, high-entropy distributions benefit more from this evaluation configuration, which suppresses low-probability “tail” generations. In contrast, STAPO learns a more intrinsically stable distribution and therefore relies less on decoding heuristics. Overall, STAPO achieves state-of-the-art performance under both configurations, demonstrating its robustness and stability.
@@ 5.3 Hyperparameter Sensitivity
(a) Sensitivity to probability threshold ($\tau_{p}$)
(b) Sensitivity to entropy threshold ($\tau_{h}$)
Figure 4: Sensitivity Analysis. Performance is reported as Acc (avg@32) on AIME24 and AIME25 benchmarks for Qwen3-1.7B base.
(a) Average Acc (avg@32) across six benchmarks with $\rho_{\mathrm{T}}$=1.0, top-p=1.0.
(b) Average Acc (avg@32) across six benchmarks with $\rho_{\mathrm{T}}$=0.7, top-p=0.9.
Figure 5: Ablation Results of Different Masking Strategies.